OpenStudy (anonymous):

Write the equation in point-slope form for a line that has a slope of 4 and passes through the point (3, 5).

OpenStudy (watchmath):

The point-slope formula is the following \(y-y_1=m(x-x_1)\) So \(y-5=4(x-3)\)